OpenStudy (anonymous):

Calculus: Find the derivative. y=4x/squareroot(2x-1)

OpenStudy (anonymous):

\[\frac{ 4x }{ \sqrt{2x-1} }\]

OpenStudy (anonymous):

u knw Quotient rule...!!

OpenStudy (anonymous):

low d high - high d low

OpenStudy (anonymous):

Hint: \[\sqrt{2x-1}= (2x-1)^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

yeah i know how to do these its just that i mite have made a mistake :\

OpenStudy (anonymous):

the answer is \[\frac{ 4x-4 }{ (2x-1)_{2}^{3} }\]

OpenStudy (anonymous):

but i got \[\frac{ 12x-4 }{ (2x-1)_{2}^{3} }\]

OpenStudy (anonymous):

when i used quotient rule :)